OpenStudy (anonymous):

Could someone please give me a hand? I don't want the answer, just someone to get me solving in the right direction. Thank you. I tried the monthly payment formula with no luck. :( Aimee and Ben are purchasing a condominium and are financing $365,000. The mortgage is a 15-year 3/1 ARM at 6.25% with a cap structure of 1.5/8. What will their payments be for the first 3 years?

jimthompson5910 (jim_thompson5910):

I'm assuming you're referring to something like this formula (shown on this page) http://www.mtgprofessor.com/formulas.htm

jimthompson5910 (jim_thompson5910):

Are you using that formula? Or something completely different?

OpenStudy (anonymous):

No, something different! More like.. bm=b ( (1+i)^nt - (1+i) ^m ----------------- (1+i)^nt-1

jimthompson5910 (jim_thompson5910):

ok that looks like the remaining balance formula

jimthompson5910 (jim_thompson5910):

bm looks like the remaining balance at month m (where m < n)

jimthompson5910 (jim_thompson5910):

which is formula #2 (or very similar to it) on that page I showed you

jimthompson5910 (jim_thompson5910):

so you're on the right track, just using the wrong formula

OpenStudy (anonymous):

Well no wonder I'm not doing it right! My math notes are very confusing. I'm taking financial algebra online. Okay, I can try that formula then. I'm just not so sure as to what all the variables that page is using stand for.

OpenStudy (anonymous):

Ah, never mind, I see it! Sorry, haha.

jimthompson5910 (jim_thompson5910):

ok on that page L = amount loaned out (or initial balance) c = interest rate per month (c = r/12.... r is the annual interest rate) n = number of months (n = 12*t .... t is the number of years)

jimthompson5910 (jim_thompson5910):

These 3 variables apply for both equations or formula, but you're only focused on the first one (the monthly payment one), so that's all you need to define

OpenStudy (anonymous):

Okay, so 365000 [ .005208 ( 1+.005208 ) ^36 / [ 1+ .005208 ) ^36 - 1 ] gave me a crazy number.... let me try it in the calculator again.

OpenStudy (anonymous):

Yeah. -363099.08. :(

OpenStudy (anonymous):

My quiz timed out :( I'll have to start over. Thank you for your help though!

jimthompson5910 (jim_thompson5910):

Sorry for some reason it just told me I had a new message

jimthompson5910 (jim_thompson5910):

btw, I think n = 12*15 = 180 and NOT n = 12*3 = 36 this is because the total life of the loan is 15 years